The sum of two numbers is 4 less than twice the larger. If the larger is decreased by three times the smaller, the result is -20. Find the numbers.
Answer by user_dude2008(1862)

x+y=2x-4
x-3y=-20

x-3y=-20
x=3y-20


x+y=2x-4
3y-20+y=2(3y-20)-4
4y-20=6y-44
2y=24
y=12

x=3y-20=3(12)-20=36-20=16

Answers: x=16, y=12